A couple is making their own wedding cake that has 3 layers. The Top layer: Diameter: 6cm Height: 3cm .... Middle Layer: Diameter 12 cm Height: 4cm...... Bottom Layer: Diameter: 20cm Height: 5cm . The couple wants to know how much to spend on frosting. If each tub of frosting costs $6.99 a tub and each tub of frosting covers 120cm^2 Figure out how much money the couple will have to spend on frosting. ( when calculating the surface area of the cylinders/layers, don't include the bottom base of each layer/cylinder because the bottom base will not get frosting. ) SA of cylinder = 2*(pi)*R*H + 2*(pi)*R^2
Answer by richwmiller(17219)

That is the wrong formula for surface area because we will never be covering the bottom of the cake and we also need to subtract the area of the layer above.
pi*r^2
3.14*6^2
pi*r^2
3.14*12^2
sa=2*pi*r*h+pi*r^2
sa=2*3.14*20*5+3.14*20^2
sa=2*3.14*12*4+3.14*12^2
sa=2*3.14*6*3+3.14*6^2
add those three and subtract
both
3.14*6^2
3.14*12^2
so we just add
sa=2*3.14*20*5+3.14*20^2
sa=2*3.14*12*4
sa=2*3.14*6*3